OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x?

hartnn (hartnn):

u get x-intercept bu putting y=0

hartnn (hartnn):

*by

OpenStudy (anonymous):

what? how would you get that?

hartnn (hartnn):

put y=0 in y + 6 = x2 − x u get 0+6=x^2-x. so x^2-x-6=0 can u solve this ?

OpenStudy (anonymous):

um...i dont think i remember..walk me?

OpenStudy (anonymous):

wait..i think i got it :P x=3 and x=-2?

OpenStudy (anonymous):

maybe? :)

hartnn (hartnn):

wow! very correct :) u are good.

OpenStudy (anonymous):

sometimes.....;)

hartnn (hartnn):

so your x-intercepts are (3,0)and(-2,0)

OpenStudy (anonymous):

yup! <3 thanks i have a few more if intrested ;)

hartnn (hartnn):

sure, just post it as a new question :) if i m busy,others can help....

OpenStudy (anonymous):

okay thanks